# LEDA: Log-Euclidean Diffeomorphic Autoencoder for Efficient Statistical Analysis of Diffeomorphisms

### v1

@@ Abstract
Image registration is a core task in computational anatomy that establishes correspondences between images. Invertible deformable registration, which computes a deformation field and handles complex, non-linear transformation, is essential for tracking anatomical variations, especially in neuroimaging applications where inter-subject differences and longitudinal changes are key. Analyzing the deformation fields is challenging due to their non-linearity, limiting statistical analysis. However, traditional approaches for analyzing deformation fields are computationally expensive, sensitive to initialization, and prone to numerical errors, especially when the deformation is far from the identity. To address these limitations, we propose the Log-Euclidean Diffeomorphic Autoencoder (LEDA), an innovative framework designed to compute the principal logarithm of deformation fields by efficiently predicting consecutive square roots. LEDA operates within a linearized latent space that adheres to the diffeomorphisms group action laws, enhancing our model’s robustness and applicability. We also introduce a loss function to enforce inverse consistency, ensuring accurate latent representations of deformation fields. Extensive experiments with the OASIS-1 dataset demonstrate the effectiveness of LEDA in accurately modeling and analyzing complex non-linear deformations while maintaining inverse consistency. Additionally, we evaluate its ability to capture and incorporate clinical variables, enhancing its relevance for clinical applications.
@@ 1 Introduction
Several methods such as Principal Geodesic Analysis (PGA) its variants [21, 47, 46], Fréchet means [33], and geodesic regression [22] have been developed to address the challenges of statistical estimation on manifolds. However, applying these methods to Fréchet Lie groups of diffeomorphisms requires significant adaptation due to the unique complexities of their structure. Adding a Riemannian metric to a Lie group to convert it into a Riemannian manifold is non-trivial, as not all Lie groups naturally possess such metrics. For example, a bi-invariant metric is a type of Riemannian metric that remains unchanged under left and right multiplication by group elements. They simplify computations, provide consistent geometric interpretations, and are particularly useful for analyzing symmetry. Compact Lie groups (e.g., $\mathrm{SO}(n)$, the group of rotations) have bi-invariant metrics; however, they are generally absent for non-compact Fréchet Lie groups such as $\mathrm{Diff}(\mathcal{M})$. This necessitates alternative meaningful metrics for non-compact groups to capture complex relationships between deformation fields in $\mathrm{Diff}(\mathcal{M})$ and ensure anatomically meaningful models.
@@
These drawbacks underscore the need for more robust and efficient approaches to estimating principal logarithms. To address these limitations, we propose Log-Euclidean Diffeomorphic Autoencoder (LEDA), an innovative framework designed to compute the principal logarithm of deformation fields by efficiently predicting the consecutive square roots of deformation fields. The framework facilitates statistical analysis within a linearized latent space that respects the group action laws of the diffeomorphism group. By appropriately mapping composition in the data space to scaling in the latent space, the framework enables the application of vector-space-based statistical methods, enhancing the robustness and applicability of statistical analysis for non-linear deformations. Furthermore, we introduce a loss function to enforce inverse consistency constraints, ensuring the latent representations accurately capture the properties of the deformation fields.
@@ 3 Background
The space of diffeomorphisms, $\mathrm{Diff}(\mathcal{M})$, forms a Lie group under composition, with the identity map, serving as the group’s identity element; i.e., ${\boldsymbol{\phi}}\circ\mathbf{id}=\mathbf{id}\circ{\boldsymbol{\phi}}={\boldsymbol{\phi}},\forall{\boldsymbol{\phi}}\in\mathrm{Diff}(\mathcal{M})$. The associated Lie algebra consists of smooth vector fields $\mathbf{v}(\mathbf{x}):\mathbb{R}^{D}\to\mathbb{R}^{D}$, which generate diffeomorphisms through their flows. These flows are described by one-parameter subgroups, $\{{\boldsymbol{\phi}}_{t}\}_{t\in\mathbb{R}}$, a continuous family of diffeomorphisms satisfying the group property: ${\boldsymbol{\phi}}_{s}\circ{\boldsymbol{\phi}}_{t}={\boldsymbol{\phi}}_{s+t}\forall s,t\in\mathbb{R},$ with ${\boldsymbol{\phi}}_{0}=\mathbf{id}$ as the identity map. The flows generated by these vector fields are governed by the ordinary differential equation (ODE) that describes how a point $\mathbf{x}$ is transported by the vector field $\mathbf{v}$ over time $t$ starting from the identity map at $t=0$:
@@
The exponential map, denoted as $\exp(\mathbf{v})$, establishes a connection between the Lie algebra and the Lie group by generating a one-parameter subgroup $\{{\boldsymbol{\phi}}_{t}\}_{t\in\mathbb{R}}$ from a vector field $\mathbf{v}$. The logarithm map $\log({\boldsymbol{\phi}})$ serves as the local inverse of the exponential map, allowing us to represent a diffeomorphism ${\boldsymbol{\phi}}$ in terms of its generating vector field. Specifically, $\log({\boldsymbol{\phi}})=\mathbf{v}$ satisfies $\exp(\mathbf{v})={\boldsymbol{\phi}}$.
Logarithm Estimation: To compute the logarithm map $\log({\boldsymbol{\phi}})$, a non-linear inverse scaling and square rooting algorithm has been proposed [6]. This approach begins by selecting a scaling factor $2^{N}$, which defines the number of successive square root operations needed to computed. These operations iteratively transform ${\boldsymbol{\phi}}$ into a version that is closer to the identity map $\mathbf{id}$, where the logarithm map is more accurately approximated. Once $N^{th}$ square root ${\boldsymbol{\phi}}^{{-2}^{N}}$ is obtained, the logarithm of ${\boldsymbol{\phi}}$ is estimated as:
@@
However, this algorithm has several drawbacks. The repeated square root operations are computationally expensive, especially for high-dimensional data or when ${\boldsymbol{\phi}}$ is far from the identity. Furthermore, the algorithm is sensitive to initialization and the iterative process can accumulate numerical errors, reducing accuracy for non-identity diffeomorphisms.
@@ 4 Methods
Figure 1: (a) Mathematical framework for diffeomorphic transformations: Shows the relationship between the Euclidean vector space (tangent space $T_{\phi}M$), the diffeomorphic group (Lie group/manifold), and the latent space ($\mathds{R}^{L}$). Transformations (${\boldsymbol{\phi}}_{AB},{\boldsymbol{\phi}}_{BA}$) and the logarithmic map ($\operatorname{log}({\boldsymbol{\phi}}_{AB})$), alongside their projections into LEDA’s latent space ($\mathbf{z}_{AB},\mathbf{z}_{BA}$), demonstrating inverse consistency and the mapping of composition in the data space to scaling in the latent space. (b) LEDA architecture
4.1 LEDA: Log-Euclidean Diffeomorphic Autoencoder
We introduce the Log-Euclidean Diffeomorphic Autoencoder (LEDA), a novel approach to efficiently estimate $N$ successive square roots of the input deformation field. Such an iterative decomposition progressively reduces the deformation field closer to the identity, offering a computationally efficient and accurate approximation of the logarithm map, even for complex, high-dimensional deformations. The LEDA architecture (Figure 1.b) includes an encoder $f_{\gamma}({\boldsymbol{\phi}})$ that maps the input ${\boldsymbol{\phi}}$ to a low-dimensional latent representation $\mathbf{z}\in\mathbb{R}^{L}$. The decoder function $g_{\theta}$ reconstructs the square roots of the deformation field, ensuring a consistent mapping between scaling in the latent space and composition in the deformation space. It predicts the $n^{th}$ root of the deformation field using scaled versions of the latent representation $\mathbf{z}$. Mathematically, this can be expressed as:
@@ 4.1 LEDA: Log-Euclidean Diffeomorphic Autoencoder
2. Inverse consistency: The estimated roots should maintain inverse consistency, i.e., ${\boldsymbol{\phi}}_{AB}({\boldsymbol{\phi}}_{BA}(\mathbf{x}))=\mathbf{x}$ for all $\mathbf{x}$. Here, $A$ and $B$ represent two coordinate spaces or imaging domains) between which the deformations occur. To facilitate this, we model LEDAas a Siamese [32] autoencoder, that processes the forward ${\boldsymbol{\phi}}_{AB}$ and inverse ${\boldsymbol{\phi}}_{BA}$ fields simultaneously.
@@ 5 Experiments
Results: Figure 2.c shows a qualitative comparison of both methods. Notably, ISS is highly sensitive to initialization values, requiring a specific initialization of $\frac{{\boldsymbol{\phi}}}{2}$. We tested multiple initializations and present results from the best-performing one. Both methods produce similar square root estimates, with higher roots approaching the identity transformation progressively (top to bottom). The log-determinant of the Jacobian reveals local deformations, where red indicates contraction and blue indicates expansion, and both methods show consistent patterns. However, ISS struggles with inverse consistency, as evidenced by the grid plots in the first row. While LEDA fully recovers the identity grid, ISS leaves residual deformations, highlighting  model’s superior robustness in maintaining inverse consistency. Moreover, estimating the square roots using ISS for a single deformation field takes 2.3 seconds (given the PyTorch implementation), whereas performing inference via the trained LEDA to estimate all the roots for a pair of deformation fields takes only 0.02 seconds. This substantial difference in computation time becomes even more pronounced when dealing with high-resolution deformation fields, underscoring the scalability and efficiency of the proposed framework.
@@ 6 Conclusion
In this work, we address the challenges of analyzing non-linear deformation fields in image registration by introducing a novel framework called Log-Euclidean Diffeomorphic Autoencoder (LEDA), designed to compute the principal logarithm of deformation fields by efficiently predicting the consecutive square roots of deformation fields. Extensive evaluations show LEDA’s effectiveness in estimating logarithm maps that capture clinically relevant anatomical variations. LEDA’s latent space can robustly link deformation fields to clinical variables, offering valuable insights into disease progression. With its efficiency and accuracy, LEDA opens avenues for efficient analysis of deformation fields, enabling more precise neuroimaging and medical applications.

### v2

# LEDA: Log-Euclidean Diffeomorphism Autoencoder for Efficient Statistical Analysis of Diffeomorphisms

@@ Abstract
Image registration is a core task in computational anatomy that establishes correspondences between images. Invertible deformable registration, which computes a deformation field and handles complex, non-linear transformation, is essential for tracking anatomical variations, especially in neuroimaging applications where inter-subject differences and longitudinal changes are key. Analyzing the deformation fields is challenging due to their non-linearity, limiting statistical analysis. However, traditional approaches for analyzing deformation fields are computationally expensive, sensitive to initialization, and prone to numerical errors, especially when the deformation is far from the identity. To address these limitations, we propose the Log-Euclidean Diffeomorphism Autoencoder (LEDA), an innovative framework designed to compute the principal logarithm of deformation fields by efficiently predicting consecutive square roots. LEDA operates within a linearized latent space that adheres to the diffeomorphisms group action laws, enhancing our model’s robustness and applicability. We also introduce a loss function to enforce inverse consistency, ensuring accurate latent representations of deformation fields. Extensive experiments with the OASIS-1 dataset demonstrate the effectiveness of LEDA in accurately modeling and analyzing complex non-linear deformations while maintaining inverse consistency. Additionally, we evaluate its ability to capture and incorporate clinical variables, enhancing its relevance for clinical applications.
@@ 1 Introduction
Several methods such as Principal Geodesic Analysis (PGA) and its variants [21, 47, 46], Fréchet means [33], and geodesic regression [22] have been developed to address the challenges of statistical estimation on manifolds. However, applying these methods to Fréchet Lie groups of diffeomorphisms requires significant adaptation due to the unique complexities of their structure. Adding a Riemannian metric to a Lie group to convert it into a Riemannian manifold is non-trivial, as not all Lie groups naturally possess such metrics. For example, a bi-invariant metric is a type of Riemannian metric that remains unchanged under left and right multiplication by group elements. They simplify computations, provide consistent geometric interpretations, and are particularly useful for analyzing symmetry. Compact Lie groups (e.g., $\mathrm{SO}(n)$, the group of rotations) have bi-invariant metrics; however, they are generally absent for non-compact Fréchet Lie groups such as $\mathrm{Diff}(\mathcal{M})$. This necessitates alternative meaningful metrics for non-compact groups to capture complex relationships between deformation fields in $\mathrm{Diff}(\mathcal{M})$ and ensure anatomically meaningful models.
@@
These drawbacks underscore the need for more robust and efficient approaches to estimating principal logarithms. To address these limitations, we propose Log-Euclidean Diffeomorphism Autoencoder (LEDA), an innovative framework designed to compute the principal logarithm of deformation fields by efficiently predicting the consecutive square roots of deformation fields. The framework facilitates statistical analysis within a linearized latent space that respects the group action laws of the diffeomorphism group. By appropriately mapping composition in the data space to scaling in the latent space, the framework enables the application of vector-space-based statistical methods, enhancing the robustness and applicability of statistical analysis for non-linear deformations. Furthermore, we introduce a loss function to enforce inverse consistency constraints, ensuring the latent representations accurately capture the properties of the deformation fields.
@@ 3 Background
The space of diffeomorphisms, $\mathrm{Diff}(\mathcal{M})$, forms a Lie group under composition, with the identity map serving as the group’s identity element; i.e., ${\boldsymbol{\phi}}\circ\mathbf{id}=\mathbf{id}\circ{\boldsymbol{\phi}}={\boldsymbol{\phi}},\forall{\boldsymbol{\phi}}\in\mathrm{Diff}(\mathcal{M})$. The associated Lie algebra consists of smooth vector fields $\mathbf{v}(\mathbf{x}):\mathbb{R}^{D}\to\mathbb{R}^{D}$, which generate diffeomorphisms through their flows. These flows are described by one-parameter subgroups, $\{{\boldsymbol{\phi}}_{t}\}_{t\in\mathbb{R}}$, a continuous family of diffeomorphisms satisfying the group property: ${\boldsymbol{\phi}}_{s}\circ{\boldsymbol{\phi}}_{t}={\boldsymbol{\phi}}_{s+t}\forall s,t\in\mathbb{R},$ with ${\boldsymbol{\phi}}_{0}=\mathbf{id}$ as the identity map. The flows generated by these vector fields are governed by the ordinary differential equation (ODE) that describes how a point $\mathbf{x}$ is transported by the vector field $\mathbf{v}$ over time $t$ starting from the identity map at $t=0$:
@@
The exponential map denoted as $\exp(\mathbf{v})$, establishes a connection between the Lie algebra and the Lie group by generating a one-parameter subgroup $\{{\boldsymbol{\phi}}_{t}\}_{t\in\mathbb{R}}$ from a vector field $\mathbf{v}$. The logarithm map $\log({\boldsymbol{\phi}})$ serves as the local inverse of the exponential map, allowing us to represent a diffeomorphism ${\boldsymbol{\phi}}$ in terms of its generating vector field. Specifically, $\log({\boldsymbol{\phi}})=\mathbf{v}$ and $\exp(\mathbf{v})={\boldsymbol{\phi}}$.
Logarithm Estimation: To compute the logarithm map $\log({\boldsymbol{\phi}})$, a non-linear inverse scaling and square rooting algorithm has been proposed [6]. This approach begins by selecting a scaling factor $2^{N}$, which defines the number of successive square root operations needed to be computed. These operations iteratively transform ${\boldsymbol{\phi}}$ into a version that is closer to the identity map $\mathbf{id}$, where the logarithm map is more accurately approximated. Once $N^{th}$ square root ${\boldsymbol{\phi}}^{{-2}^{N}}$ is obtained, the logarithm of ${\boldsymbol{\phi}}$ is estimated as:
@@
However, this algorithm has several drawbacks. The repeated square root operations are computationally expensive, especially for high-dimensional data or when ${\boldsymbol{\phi}}$ is far from the identity. Furthermore, the algorithm is sensitive to initialization, and the iterative process can accumulate numerical errors, reducing the accuracy of non-identity diffeomorphisms.
@@ 4 Methods
Figure 1: (a) Mathematical framework for diffeomorphic transformations: Shows the relationship between the Euclidean vector space (tangent space $T_{\phi}M$), the diffeomorphic group (Lie group/manifold), and the latent space ($\mathds{R}^{L}$). Here, A and B represent two coordinate spaces or imaging domains between which the deformations occurs. Transformations (${\boldsymbol{\phi}}_{AB},{\boldsymbol{\phi}}_{BA}$) and the logarithmic map ($\operatorname{log}({\boldsymbol{\phi}}_{AB})$), alongside their projections into LEDA’s latent space ($\mathbf{z}_{AB},\mathbf{z}_{BA}$), demonstrating inverse consistency and the mapping of composition in the data space to scaling in the latent space. (b) LEDA architecture
4.1 LEDA: Log-Euclidean Diffeomorphism Autoencoder
We introduce the Log-Euclidean Diffeomorphism Autoencoder (LEDA), a novel approach to efficiently estimate $N$ successive square roots of the input deformation field. Such an iterative decomposition progressively reduces the deformation field closer to the identity, offering a computationally efficient and accurate approximation of the logarithm map, even for complex, high-dimensional deformations. The LEDA architecture (Figure 1.b) includes an encoder $f_{\gamma}({\boldsymbol{\phi}})$ that maps the input ${\boldsymbol{\phi}}$ to a low-dimensional latent representation $\mathbf{z}\in\mathbb{R}^{L}$. The decoder function $g_{\theta}$ reconstructs the square roots of the deformation field, ensuring a consistent mapping between scaling in the latent space and composition in the deformation space. It predicts the $n^{th}$ root of the deformation field using scaled versions of the latent representation $\mathbf{z}$. Mathematically, this can be expressed as:
@@ 4.1 LEDA: Log-Euclidean Diffeomorphism Autoencoder
2. Inverse consistency: The estimated roots should maintain inverse consistency, i.e., ${\boldsymbol{\phi}}_{AB}({\boldsymbol{\phi}}_{BA}(\mathbf{x}))=\mathbf{x}$ for all $\mathbf{x}$. Here, $A$ and $B$ represent two coordinate spaces or imaging domains between which the deformations occur. To facilitate this, we model LEDA as a Siamese [32] autoencoder, that processes the forward ${\boldsymbol{\phi}}_{AB}$ and inverse ${\boldsymbol{\phi}}_{BA}$ fields simultaneously.
@@ 5 Experiments
Results: Figure 2.c shows a qualitative comparison of both methods. Notably, ISS is highly sensitive to initialization values, requiring a specific initialization of $\frac{{\boldsymbol{\phi}}}{2}$. We tested multiple initializations and present results from the best-performing one. Both methods produce similar square root estimates, with higher roots approaching the identity transformation progressively (top to bottom). The log-determinant of the Jacobian reveals local deformations, where red indicates contraction and blue indicates expansion, and both methods show consistent patterns. However, ISS struggles with inverse consistency, as evidenced by the grid plots in the first row. While LEDA fully recovers the identity grid, ISS leaves residual deformations, highlighting LEDA’s superior robustness in maintaining inverse consistency. Moreover, estimating the square roots using ISS for a single deformation field takes 2.3 seconds (given the PyTorch implementation), whereas performing inference via the trained LEDA to estimate all the roots for a pair of deformation fields takes only 0.02 seconds. This substantial difference in computation time becomes even more pronounced when dealing with high-resolution deformation fields, underscoring the scalability and efficiency of the proposed framework.
@@ 6 Conclusion
In this work, we address the challenges of analyzing non-linear deformation fields in image registration by introducing a novel framework called Log-Euclidean Diffeomorphism Autoencoder (LEDA), designed to compute the principal logarithm of deformation fields by efficiently predicting the consecutive square roots of deformation fields. Extensive evaluations show LEDA’s effectiveness in estimating logarithm maps that capture clinically relevant anatomical variations. LEDA’s latent space can robustly link deformation fields to clinical variables, offering valuable insights into disease progression. With its efficiency and accuracy, LEDA opens avenues for efficient analysis of deformation fields, enabling more precise neuroimaging and medical applications.
Acknowledgement
This work was supported by the National Institutes of Health under grant numbers NIH-R01CA259686 (Sarang Joshi) and NIH-R01DE032366 (Shireen Elhabian).